Pearson Motors has a target capital structure of 40% debt and 60% common equity, with no preferred stock. The yield to maturity on the company's outstanding bonds is 12%, and its tax rate is 25%. Pearson's CFO estimates that the company's WACC is 10.40%. What is Pearson's cost of common equity?
Answer #1

After-tax cost of debt=yield to maturity*(1-tax rate)

=12*(1-0.25)=9%

WACC=Respective costs*Respective weight

10.4=(9*0.4)+(0.6*Cost of common equity)

Cost of common equity=(10.4-3.6)/0.6

=11.33%(Approx).